OpenStudy (anonymous):

Acceleration (a) of a moving object is directly proportional to its velocity (v). If velocity is 4 when acceleration is 10, what is acceleration when velocity is 27.5

OpenStudy (mayankdevnani):

\[\huge \bf acceleration(a)=k \times velocity(v)\] where k is constant. If v=4 and a=10, \[\huge \bf k=\frac{a}{v}=\frac{10}{4}=\frac{5}{2}\]

OpenStudy (mayankdevnani):

If v=27.5 and a=? \[\huge \bf k=\frac{a}{v}\] \[\huge \bf a=k \times v\] so, \[\huge \bf a=\frac{5}{2} \times 27.5\]

OpenStudy (mayankdevnani):

Can you solve for a? @balokkk

OpenStudy (anonymous):

good boi @mayankdevnani =]

OpenStudy (mayankdevnani):

thank you :) @genius12

OpenStudy (anonymous):

68.75? @mayankdevnani

OpenStudy (mayankdevnani):

yeah,you are correct

OpenStudy (anonymous):

thank youu

OpenStudy (mayankdevnani):

welcome :)